Given the information below, provide the Adjusting Journal Entries for the following business transactions by TechCo:

  1. On Jan 1, 2020, TechCo's company provided education services to customers in the amount of $15,500. The company will collect the payment later.

  2. TechCo's company recognized February’s insurance cost. On Jan 1, the company had paid $42,000 for 8 months coverage from Jan 1 to Aug 31.

  3. TechCo's company provided another education service on March 15, 2020, totaling $20,500 to customers that had paid for services in advance.

  4. On March 31, 2020, TechCo's company recognized its interest owed on a mortgage payable for $1,650.

  5. On April 30, 2020, TechCo's company recognized the depreciation of a factory machine $2,000 for the current month.

  6. In May, TechCo's company counted the office supplies that are used during the month. At the beginning of the month, the total of office supplies on hand was $3,000, while $2,400 of supplies were left at the end of May. During the month, $25,000 of additional supplies were purchased.

  7. Employees at TechCo's company work 5 days every week and the weekly wages total $60,000. On June 17, 2020, employees earned accrued wages for the last 3 days of the month.

  8. TechCo company earns a 6% annual interest on its $100,000 of investments. Interest is paid every six months on June 30 and December 31. If monthly financial statements are prepared, what adjusting entry should be made on Jan 31, 2021?

Please provide Adjusting Journal Entries for the above transactions.
Answer #1
Account Titles Debit Credit
a Accounts Receivable $          15,500
        Education Service Revenue     $       15,500
b Insurance Expense $            5,250 =42000/8
        Prepaid Insurance $         5,250
c Unearned Revenue $          20,500
        Education Service Revenue $       20,500
d Interest Expense $            1,650
        Interest Payable $         1,650
e Depreciation Expense $            2,000
        Accumulated Depreciation $         2,000
f Supplies Expense $          25,600 =3000+25000-2400
       Office Supplies $       25,600
g Wages Expense $          36,000 =60000/5*3
       Wages Payable $       36,000
h Interest Expense $                500 =100000*6%*1/12
       Interest Payable $             500

If you have any query, kindly comment with your query and please mark thumbs up.